OpenStudy (anonymous):

a square has a perimeter 40 inches. What would be the perimeter of an equilateral triangle whose sides each measure the same length as the side of the square?

OpenStudy (anonymous):

The square sides = 40 / 4 = 10 What is the perimeter of the equilateral triangle?

OpenStudy (anonymous):

that's what i need to know the perimeter

OpenStudy (jjuden):

30

OpenStudy (anonymous):

@talita Re-read this "What would be the perimeter of an equilateral triangle whose sides each measure the same length as the side of the square?" I gave you the sides of the square where I stated The square sides = 40 / 4 = 10 The square sides are 10. A square has 4 sides. A triangle has 3 sides. An equilateral triangle has 3 sides that are all the same. 10 + 10 + 10= 30 30 is the perimeter of the triangle.

OpenStudy (anonymous):

You are correct @jjuden

OpenStudy (jjuden):

@MDoodler i know lol

OpenStudy (anonymous):

thank you MDoodler